ID: A 18. as points) A standing wave is set up in a segments driven by a (a) What is the wavelength of wave? a. 10 cm the 20 cm c. 40 ing fixed at both ends. The string vibrates in s distinct 80 cm harmonic (b) what is the fundamental requency a. 48 Hz b. 122 (20) ng? d. 86 Hz e, 240 Hz SL (c) If the mass per unit length of the string is 2.5 g/m, what is the tension in the a. 120N 23 d. 460 N 100 c. 23 000 N string? S a L N M L d00 cm 2 m 2.5 gl if I had the fundamental frequent, T could solve for tension d the strin Part B & C please.
Answer #1


f =120 Hz, L = 200 cm

lamda = 2L/5 = 2*200/5 = 80 cm

(b) f5 = 5*fo

120 = 5*fo

fo = 24 Hz

(c) fo = (1/2L)(T/u)^0.5

u = 2.5 g/m

24 = (1/(2*2))(T/(2.5*0.001))^0.5

T = 23 N